OpenStudy (kailee1423):

Divide -3x3 - 2x2 - x - 2 by x – 2.

OpenStudy (kailee1423):

Can someone please help?

OpenStudy (kailee1423):

Wait, it didn't post right. It's Divide -3x^3 - 2x^2 - x - 2 by x – 2.

OpenStudy (evoker):

I would use long division

OpenStudy (kailee1423):

I suck at long division.... Can you explain it?

OpenStudy (evoker):



OpenStudy (faiqraees):

@Evoker Won't synthetic division would be the most neat and easy choice here?

OpenStudy (evoker):

Sure ok there is the setup problem

OpenStudy (evoker):

You could use that, it could be easy, but long division is more generally useful.

OpenStudy (kailee1423):

Whichever is easiest to understand. Lol

OpenStudy (evoker):

For understanding I like long division, so anyways start with the two front pieces

OpenStudy (evoker):

so -3x^3/x is what?

OpenStudy (evoker):

In otherwords the front term inside the division region divided by the front term of the divisor.

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Kailee1423 Whichever is easiest to understand. Lol \(\color{#0cbb34}{\text{End of Quote}}\) The easiest to understand is the synthetic one but the one @Evoker is providing is though a little hard but will help you later on too.

OpenStudy (evoker):

Anyways I have to head out for now , I'll let FaiqRaess take over.

OpenStudy (kailee1423):

Ok thanks anyways!

OpenStudy (faiqraees):

@Kailee1423 Which do you prefer, synthetic or long division?

OpenStudy (kailee1423):

I don't understand either of them so it doesn't matter to me

OpenStudy (faiqraees):

So I am gonna start with synthetic. Fine by you?

OpenStudy (kailee1423):

Sure:)

OpenStudy (faiqraees):

Okay so first thing. Can you convert a factor into root? (Hint (x-2) is a factored form and x=2 is a root)

OpenStudy (kailee1423):

No, I'm just barely passing my math class. All of it is confusing

OpenStudy (faiqraees):

Okay so if something is divisible by a monomial or polynomial, the divisor is considered as a factor. Like if x²-2x+1 is divisible by (x-1). (x-1) is considered a factor. Clear on that?

OpenStudy (kailee1423):

Yes

OpenStudy (faiqraees):

if x-1 is a factor. You can easily find out a root. The procedure is to equate the factor to 0 (x-1)=0 x-1=0 x=1 Hence x=1 is a root. Clear? Any confusions so far?

OpenStudy (kailee1423):

Kinda....

OpenStudy (faiqraees):

Go on elaborate your confusion.

OpenStudy (kailee1423):

Wait... I get it

OpenStudy (faiqraees):

Should we proceed?

OpenStudy (kailee1423):

Yes

OpenStudy (faiqraees):

So first we are going to convert our factor into a root. Can you convert (x-2) into a root form? Hint:-(Just equate it to 0)

OpenStudy (kailee1423):

(X-2)=0 X=2

OpenStudy (faiqraees):

Perfect. Now we're going to divide our polynomial with the root.  This is our setup. Clear?

OpenStudy (kailee1423):

Ok

OpenStudy (faiqraees):

Btw are you sure the question is correct?

OpenStudy (kailee1423):

Yes. Why?

OpenStudy (faiqraees):

Oh then apologies for inconvience. The question can't be solved by synthetic division. We have to apply long division. Apologies for wasting your time. But if you want I can explain the long division..

OpenStudy (kailee1423):

It's ok! I need to hurry up and get the assignment turned in. Does long division take a while?

OpenStudy (faiqraees):

Yes a little, how about I provide you an article or something, which would speed the process a little?

OpenStudy (kailee1423):

That would be great:)

OpenStudy (faiqraees):

http://www.purplemath.com/modules/polydiv2.htm

OpenStudy (kailee1423):

Thank you!

OpenStudy (faiqraees):

No problem